# Fair Wasserstein Coresets

### v3

@@ 1 Introduction
In the last decade, the rapid pace of technological advancement has provided the ability of collecting, storing and processing massive amounts of data from multiple sources (Sagiroglu & Sinanc, 2013). As the volume of data continues to surge, it often surpasses both the available computational resources as well as the capacity of machine learning algorithms. In response to this limitation, dataset distillation approaches aim to reduce the amount of data by creating a smaller, yet representative, set of samples; see (Yu et al., 2024; Lei & Tao, 2024) for comprehensive reviews on the topic. Among those approaches, coresets provide a weighted subset of the original data that achieve similar performance to the original dataset in (usually) a specific machine learning task, such as clustering (Har-Peled & Mazumdar, 2004; Feldman, 2020), Bayesian inference (Campbell & Broderick, 2018), online learning (Borsos et al., 2020) and classification (Coleman et al., 2020), among others.
In tandem with these developments, the adoption of machine learning techniques has seen a surge in multiple decision-making processes that affect society at large (Sloane et al., 2022; Zhang et al., 2022). This proliferation of machine learning applications has highlighted the need to mitigate inherent biases in the data, as these biases can significantly impact the equity of machine learning models and their decisions (Chouldechova, 2017). Among many definitions of algorithmic fairness, demographic parity is one of the most prominently used metric (Hort et al., 2023), enforcing the distribution of an outcome of a machine learning model to not differ dramatically across different subgroups in the data.
Current methodologies for generating a smaller set of fair representative samples focus on the local characteristics of these samples with respect to the original dataset. For instance, Chierichetti et al. (2017); Huang et al. (2019); Backurs et al. (2019); Ghadiri et al. (2021) obtain representative points by clustering while enforcing each cluster to include the same proportion of points from each subgroup in the original dataset. In another line of work, Jung et al. (2019); Mahabadi & Vakilian (2020); Negahbani & Chakrabarty (2021); Vakilian & Yalciner (2022); Chhaya et al. (2022) create representative points by ensuring that points in the original dataset each have at least one representative point within a given distance in the feature space. While these methods can successfully reduce clustering cost and ensure a more evenly spread-out distribution of representative points in the feature space, it is unclear whether such representative samples can positively affect performance or discrimination reduction in downstream learning processes. As the induced distribution of the representative points might be far away from the original dataset distribution, downstream machine learning algorithm might lose significant performance due to this distribution shift, without necessarily reducing biases in the original data (as we also demonstrate empirically in our experiments).
@@
    we develop an efficient majority minimization algorithm (Ortega & Rheinboldt, 2000; Lange, 2016) to solve the reformulated problem (Section 4). We analyze theoretical properties of our proposed algorithm and FWC (Section 5) and show that, in the absence of fairness constraints, our algorithm reduces to an equivalent version of Lloyd’s algorithm for k-means and k-medians clustering, extending its applicability beyond fairness applications (Section 6);
@@
    we empirically validate the scalability and effectiveness of FWC by providing experiments on both synthetic and real datasets (Section 7). In downstream learning tasks, FWC result in competitive fairness-utility tradeoffs against current approaches, even when we enhance the fairness of existing approaches using fair pre-processing techniques, with an average disparity reduction of $53\%$ and $18\%$, respectively. In addition, we show FWC can correct biases in large language models when passing coresets as examples to the LLM, reducing downstream disparities by $75\%$ with GPT-3.5 (OpenAI, 2022) by $35\%$ with GPT-4 (Achiam et al., 2023). Finally, we show FWC can improve downstream fairness-utility tradeoffs in downstream models when added to the training data (via data augmentation, see Appendix C.2).
@@ 2 Background and Related Work
with $\Pi(p_{1},p_{2})$ indicating the set of all joint probability distributions over the product space $\mathcal{X}\times\mathcal{X}$ with marginals equal to $(p_{1},p_{2})$ (Kantorovitch, 1958). The operator $c(x,y)$ represents the “cost” of moving probability mass from $x$ and $y$, and reduces to a matrix $C$ if the underlying metric space ${\cal X}$ is discrete.
The Wasserstein distance has several connections with downstream learning processes and coresets. Firstly, the higher the Wasserstein distance between the weighted representative samples $p_{\widehat{Z};\theta}$ and the original dataset $p_{Z;e}$, the higher the distribution shift, the more degradation we might expect in terms of downstream learning performance (Quiñonero-Candela et al., 2022). Secondly, the Wasserstein distance between two probability distributions can also be used to bound the deviation of functions applied to samples from such distributions. Define the following deviation:
@@
When ${\cal F}$ is the class of Lipschitz-continuous functions with Lipschitz constant equal or less than $1$, the deviation $d(p_{\widehat{Z};\theta},p_{Z;e})$ is equal to 1-Wasserstein distance $\mathcal{W}_{1}(p_{\widehat{Z};\theta},p_{Z;e})$ (Santambrogio, 2015; Villani et al., 2009). The connection with downstream learning processes is immediate when considering Lipschitz continuous classifiers, such as logistic regression or Lipschitz-constrained neural networks (Anil et al., 2019). For other classifiers, we note that the 1-Wasserstein distance still bounds the downstream discrepancy: in Proposition 2.1 we show that the 1-Wasserstein distance bounds the downstream discrepancy for ReLu-activated multilayer perceptrons (MLPs), which we use in our experiments on real datasets (Section 7).
@@
Let $g_{\psi}\in\mathcal{G}^{K}$ be the class of $K$-layer multilayer perceptrons with ReLu activations. Then, the downstream discrepancy in downstream performance of $g_{\psi}$ applied to samples from $p_{\widehat{Z};\theta}$ and $p_{Z;e}$ is bounded by the 1-Wasserstein distance :
$d(p_{\widehat{Z};\theta},p_{Z;e})=\sup_{g_{\psi}\in\mathcal{G^{K}}}\left|\mathbb{E}_{z\sim p_{\widehat{Z};\theta}}g_{\psi}(z)-\mathbb{E}_{z\sim p_{Z;e}}g_{\psi}(z)\right|\leq L_{k}\mathcal{W}_{1}(p_{\widehat{Z};\theta},p_{Z;e}),$ (2)
where $L_{k}$ is the MLP Lipschitz constant upper bound defined in Virmaux & Scaman (2018, Section 6.1, Equation (8)).
We point out that minimizing the Wasserstein distance, hence bounding the downstream performance as in Equation (2), is equivalent to the definition of a measure coresets proposed by Claici et al. (2018). The Wasserstein distance is also connected to coresets in the sense of the best discrete approximation of a continuous distributions. Considering a feature space endowed with a continuous distribution, minimizing the $p$-Wasserstein distance across all the distributions of size $m$ is topologically equivalent to identify the samples of size $m$ that provide the best Voronoi tessellation of the space in $L_{p}$ sense (Pagès, 2015; Liu & Pagès, 2020). Although other definitions of coresets using Wasserstein distance have been proposed in the literature, they require either to solve the underlying optimal transport problems or the knowledge of the downstream classifier and loss function (Claici et al., 2018; Mirzasoleiman et al., 2020; Zhao et al., 2021; Loo et al., 2022). FWC is agnostic to any downstream model or loss function, and uses an efficient implementation that does not actually incur in the usual high cost connected to optimal transport. By adapting the approach proposed by Xiong et al. (2024), we solve an equivalent re-formulated linear optimization problem, which is more computationally tractable than classic approaches such as the simplex or the interior point method (see Section 4.1).
@@
Also known as statistical parity, demographic parity (DP) imposes the decision outcome and protected attributes to be independent (Dwork et al., 2012). Using a credit card approval decision example, demographic parity enforces an automatic decision process to approve similar proportion of applicants across different race demographics. DP is one of the most extensively analyzed fairness criterion; we refer the reader to Hort et al. (2023) for a review. In this work, we use the demographic parity definition by Calmon et al. (2017), which enforces the ratio between the conditional distribution of the decision outcome across each subgroups $p(y|D=d)$ and the marginal distribution of the decision outcome $p(y)$ to be close to $1$ in a given dataset. We refer to the “empirical version” of demographic parity to indicate that the conditional and marginal distributions are quantities estimated from the data. Note that the demographic parity definition we adopt enforces a condition on the weighted average of the conditional distributions across groups, which is different from a more recent approach of using Wasserstein distance, and specifically Wasserstein barycenters, to enforce demographic parity (Gordaliza et al., 2019; Jiang et al., 2020b; Gaucher et al., 2023; Xu & Strohmer, 2023).
@@ 3 FWC: Fair Wasserstein Coresets
Given a dataset $\{Z_{i}\}_{i=1}^{n}$, our goal is to find a set of samples $\{\widehat{Z}_{j}\}_{j=1}^{m}$ and weights $\{\theta_{j}\}_{j=1}^{m}$ such that $m\ll n$ and that the Wasserstein distance between $p_{Z;e}$ and $p_{\widehat{Z};\theta}$ is as small as possible. In addition, we use the fairness constraints proposed by Calmon et al. (2017) to control the demographic parity violation for the $p_{\widehat{Z};\theta}$ distribution. Let $p_{\widehat{Z};\theta}(y|d)$ indicate the conditional distribution $p_{\widehat{Z};\theta}(\widehat{Y}=y|\widehat{D}=d)$. Imposing a constraint on the demographic disparity violation then reduces to requiring the conditional distribution under the weights $\{\theta_{i}\}_{i\in[n]}$ to be close to a target distribution $p_{Y_{T}}$ across all possible values of the protected attributes $D$,
@@
where $\epsilon$ is a parameter that determines the maximum fairness violation, and $J(\cdot,\cdot)$ is the probability ratio between distributions as defined in Calmon et al. (2017).
@@
where $\Delta_{m}$ indicates the set of valid weights $\{\theta\in\mathbb{R}^{m}_{+}:\sum_{i=1}^{m}\theta_{i}=m\}$. Note that the optimization problem in (4) shares some similarities with the optimization problem in Xiong et al. (2024) in using the Wasserstein distance as a distance metric between distributions and using (3) to enforce demographic parity. However, Xiong et al. (2024) only provide sample-level integer weights for the original dataset and do not generate any new samples, while our approach provides a separate set of samples $\{\widehat{Z}_{j}\}_{j=1}^{m}$ with associated real-valued weights $\{\theta_{j}\}_{j=1}^{m}$, with $m\ll n$.
@@
For any $m>0$, the best fair Wasserstein coreset formed by $m$ data points $\{\widehat{Z}_{i}:i\in[m]\}$ is no better than the best fair Wasserstein coreset formed by $m|{\cal D}||{\cal Y}|$ data points $\{(d,X_{i},y)_{i}:i\in[m],d\in{\cal D},y\in{\cal Y}\}$.
@@
Following Xiong et al. (2024), the fairness constraint in Equation (3) can be expressed as $2|{\cal Y}||{\cal D}|$ linear constraints on the weights $\theta$, as the disparity reduces to the following for all $d\in\mathcal{D},y\in{\cal Y}$:
@@
When keeping the samples $\widehat{X}$ fixed, we can follow Peyré et al. (2019) to derive an equivalent formulation of the Wasserstein distance in the objective as a linear program with $mn$ variables. By indicating the transportation cost matrix $C(\widehat{X})$, we define its components as follows,
@@
s.t. $P\mathbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},\ P\geq 0,\ AP^{\top}\mathbf{1}_{n}\geq 0\ .$
@@
s.t. $P\mathbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},\ P\geq 0,\ AP^{\top}\mathbf{1}_{n}\geq 0$
@@ 4 Majority Minimization for Solving the Reformulated Problem
Majority minimization aims at solving nonconvex optimization problems, and refers to the process of defining a convex surrogate function that upper bounds the nonconvex objective function, so that optimizing the surrogate function improves the objective function (Ortega & Rheinboldt, 2000; Lange, 2016). As the algorithm proceeds, the surrogate function also updates accordingly, which ensures the value of the original objective function keeps decreasing. Following this framework, we define the surrogate function $g(\cdot;\widehat{X}^{k})$ as follows for the $k$-th iterate $\widehat{X}^{k}\in{\cal X}^{m}$:
@@ 4.1 Updating the Surrogate Function (Line 4)
To update the surrogate function, we need to solve problem (8), which is a large-scale linear program. Rather than solving the computationally prohibitive dual problem we solve a lower-dimensional dual problem by using a variant of the FairWASP algorithm proposed by Xiong et al. (2024). We adapt FairWASP for cases where $m\neq n$ to find the solution of (8) via applying the cutting plane methods on the Lagrangian dual problems with reduced dimension [^2] As opposed to the scenario where $m=n$, which was tackled by Xiong et al. (2024).. We choose FairWASP over established commercial solvers due to its computational complexity being lower than other state of the art approaches such as interior-point or simplex method; see Lemma A.2 in Appendix A for more details.

### v4

@@ 1 Introduction
In the last decade, the rapid pace of technological advancement has provided the ability of collecting, storing and processing massive amounts of data from multiple sources [65]. As the volume of data continues to surge, it often surpasses both the available computational resources as well as the capacity of machine learning algorithms. In response to this limitation, dataset distillation approaches aim to reduce the amount of data by creating a smaller, yet representative, set of samples; see [80, 40] for comprehensive reviews on the topic. Among those approaches, coresets provide a weighted subset of the original data that achieve similar performance to the original dataset in (usually) a specific machine learning task, such as clustering [26, 21], Bayesian inference [11], online learning [9] and classification [16], among others.
In tandem with these developments, the adoption of machine learning techniques has seen a surge in multiple decision-making processes that affect society at large [69, 81]. This proliferation of machine learning applications has highlighted the need to mitigate inherent biases in the data, as these biases can significantly impact the equity of machine learning models and their decisions [14]. Among many definitions of algorithmic fairness, demographic parity is one of the most prominently used metric [29], enforcing the distribution of an outcome of a machine learning model to not differ dramatically across different subgroups in the data.
Current methodologies for generating a smaller set of fair representative samples focus on the local characteristics of these samples with respect to the original dataset. For instance, [13, 30, 4, 24] obtain representative points by clustering while enforcing each cluster to include the same proportion of points from each subgroup in the original dataset. In another line of work, [33, 44, 53, 72, 12] create representative points by ensuring that points in the original dataset each have at least one representative point within a given distance in the feature space. While these methods can successfully reduce clustering cost and ensure a more evenly spread-out distribution of representative points in the feature space, it is unclear whether such representative samples can positively affect performance or discrimination reduction in downstream learning processes. As the induced distribution of the representative points might be far away from the original dataset distribution, downstream machine learning algorithm might lose significant performance due to this distribution shift, without necessarily reducing biases in the original data (as we also demonstrate empirically in our experiments).
@@
    we develop an efficient majority minimization algorithm [56, 39] to solve the reformulated problem (Section 4). We analyze theoretical properties of our proposed algorithm and FWC (Section 5) and show that, in the absence of fairness constraints, our algorithm reduces to an equivalent version of Lloyd’s algorithm for k-means and k-medians clustering, extending its applicability beyond fairness applications (Section 6);
@@
    we empirically validate the scalability and effectiveness of FWC by providing experiments on both synthetic and real datasets (Section 7). In downstream learning tasks, FWC result in competitive fairness-utility tradeoffs against current approaches, even when we enhance the fairness of existing approaches using fair pre-processing techniques, with an average disparity reduction of $53\%$ and $18\%$, respectively. In addition, we show FWC can correct biases in large language models when passing coresets as examples to the LLM, reducing downstream disparities by $75\%$ with GPT-3.5 [55] by $35\%$ with GPT-4 [1]. Finally, we show FWC can improve downstream fairness-utility tradeoffs in downstream models when added to the training data (via data augmentation, see Appendix C.2).
@@ 2 Background and Related Work
with $\Pi(p_{1},p_{2})$ indicating the set of all joint probability distributions over the product space $\mathcal{X}\times\mathcal{X}$ with marginals equal to $(p_{1},p_{2})$ [35]. The operator $c(x,y)$ represents the “cost” of moving probability mass from $x$ and $y$, and reduces to a matrix $C$ if the underlying metric space ${\cal X}$ is discrete.
The Wasserstein distance has several connections with downstream learning processes and coresets. Firstly, the higher the Wasserstein distance between the weighted representative samples $p_{\widehat{Z};\theta}$ and the original dataset $p_{Z;e}$, the higher the distribution shift, the more degradation we might expect in terms of downstream learning performance [62]. Secondly, the Wasserstein distance between two probability distributions can also be used to bound the deviation of functions applied to samples from such distributions. Define the following deviation:
@@
When ${\cal F}$ is the class of Lipschitz-continuous functions with Lipschitz constant equal or less than $1$, the deviation $d(p_{\widehat{Z};\theta},p_{Z;e})$ is equal to 1-Wasserstein distance $\mathcal{W}_{1}(p_{\widehat{Z};\theta},p_{Z;e})$ [66, 74]. The connection with learning processes and the downstream deviation $d(p_{(\widehat{X},\widehat{D});\theta},p_{(X,D);e})$ is immediate when considering Lipschitz continuous classifiers with Lipschitz constant less than 1 (such as logistic regression or Lipschitz-constrained neural networks [2]) [^1] In such cases, $d(p_{(\widehat{X},\widehat{D});\theta},p_{(X,D);e})=\mathcal{W}_{1}(p_{(\widehat{X},\widehat{D});\theta},p_{(X,D);e})\leq\mathcal{W}_{1}(p_{\widehat{Z};\theta},p_{Z;e})$, see Lemma B.1.. For other classifiers, we note that the 1-Wasserstein distance still bounds the downstream discrepancy: in Proposition 2.1 we show that the 1-Wasserstein distance bounds the downstream discrepancy for ReLu-activated multilayer perceptrons (MLPs), which we use in our experiments on real datasets (Section 7).
@@
Let $g_{\psi}\in\mathcal{G}^{K}$ be the class of $K$-layer multilayer perceptrons with ReLu activations. Then, the downstream discrepancy in downstream performance of $g_{\psi}$ applied to samples from $p_{(\widehat{X},\widehat{D});\theta}$ and $p_{(X,D);e}$ is bounded by the 1-Wasserstein distance :
$d(p_{(\widehat{X},\widehat{D});\theta},$ $p_{(X,D);e})=$
$\sup_{g_{\psi}\in\mathcal{G^{K}}}$ $\left|\mathbb{E}_{(x,d)\sim p_{(\widehat{X},\widehat{D});\theta}}g_{\psi}(x,d)-\mathbb{E}_{(x,d)\sim p_{(X,D);e}}g_{\psi}(x,d)\right|\leq L_{k}\mathcal{W}_{1}(p_{\widehat{Z};\theta},p_{Z;e}),$ (2)
where $L_{k}$ is the MLP Lipschitz constant upper bound defined in [75, Section 6.1, Equation (8)].
We point out that minimizing the Wasserstein distance, hence bounding the downstream performance as in Equation (2), is equivalent to the definition of a measure coresets proposed by [15]. The Wasserstein distance is also connected to coresets in the sense of the best discrete approximation of a continuous distributions. Considering a feature space endowed with a continuous distribution, minimizing the $p$-Wasserstein distance across all the distributions of size $m$ is topologically equivalent to identify the samples of size $m$ that provide the best Voronoi tessellation of the space in $L_{p}$ sense [57, 41]. Although other definitions of coresets using Wasserstein distance have been proposed in the literature, they require either to solve the underlying optimal transport problems or the knowledge of the downstream classifier and loss function [15, 46, 82, 43]. FWC is agnostic to any downstream model or loss function, and uses an efficient implementation that does not actually incur in the usual high cost connected to optimal transport. By adapting the approach proposed by [77], we solve an equivalent re-formulated linear optimization problem, which is more computationally tractable than classic approaches such as the simplex or the interior point method (see Section 4.1).
@@
Also known as statistical parity, demographic parity (DP) imposes the decision outcome and protected attributes to be independent [18]. Using a credit card approval decision example, demographic parity enforces an automatic decision process to approve similar proportion of applicants across different race demographics. DP is one of the most extensively analyzed fairness criterion; we refer the reader to [29] for a review. In this work, we use the demographic parity definition by [10], which enforces the ratio between the conditional distribution of the decision outcome across each subgroups $p(y|D=d)$ and the marginal distribution of the decision outcome $p(y)$ to be close to $1$ in a given dataset. We refer to the “empirical version” of demographic parity to indicate that the conditional and marginal distributions are quantities estimated from the data. Note that the demographic parity definition we adopt enforces a condition on the weighted average of the conditional distributions across groups, which is different from a more recent approach of using Wasserstein distance, and specifically Wasserstein barycenters, to enforce demographic parity [25, 32, 23, 78].
@@ 3 FWC: Fair Wasserstein Coresets
Given a dataset $\{Z_{i}\}_{i=1}^{n}$, our goal is to find a set of samples $\{\widehat{Z}_{j}\}_{j=1}^{m}$ and weights $\{\theta_{j}\}_{j=1}^{m}$ such that $m\ll n$ and that the Wasserstein distance between $p_{Z;e}$ and $p_{\widehat{Z};\theta}$ is as small as possible. In addition, we use the fairness constraints proposed by [10] to control the demographic parity violation for the $p_{\widehat{Z};\theta}$ distribution. Let $p_{\widehat{Z};\theta}(y|d)$ indicate the conditional distribution $p_{\widehat{Z};\theta}(\widehat{Y}=y|\widehat{D}=d)$. Imposing a constraint on the demographic disparity violation then reduces to requiring the conditional distribution under the weights $\{\theta_{i}\}_{i\in[n]}$ to be close to a target distribution $p_{Y_{T}}$ across all possible values of the protected attributes $D$,
@@
where $\epsilon$ is a parameter that determines the maximum fairness violation, and $J(\cdot,\cdot)$ is the probability ratio between distributions as defined in [10].
@@
where $\Delta_{m}$ indicates the set of valid weights $\{\theta\in\mathbb{R}^{m}_{+}:\sum_{i=1}^{m}\theta_{i}=m\}$. Note that the optimization problem in (4) shares some similarities with the optimization problem in [77] in using the Wasserstein distance as a distance metric between distributions and using (3) to enforce demographic parity. However, [77] only provide sample-level integer weights for the original dataset and do not generate any new samples, while our approach provides a separate set of samples $\{\widehat{Z}_{j}\}_{j=1}^{m}$ with associated real-valued weights $\{\theta_{j}\}_{j=1}^{m}$, with $m\ll n$.
@@
For any $m>0$, the best fair Wasserstein coreset formed by $m$ data points $\{\widehat{Z}_{i}:i\in[m]\}$ is no better (i.e., the optimal Wasserstein distance value is no lower) than the best fair Wasserstein coreset formed by $m|{\cal D}||{\cal Y}|$ data points $\{(d,X_{i},y)_{i}:i\in[m],d\in{\cal D},y\in{\cal Y}\}$.
@@
Following [77], the fairness constraint in Equation (3) can be expressed as $2|{\cal Y}||{\cal D}|$ linear constraints on the weights $\theta$, as the disparity reduces to the following for all $d\in\mathcal{D},y\in{\cal Y}$:
@@
When keeping the samples $\widehat{X}$ fixed, we can follow [60] to derive an equivalent formulation of the Wasserstein distance in the objective as a linear program with $mn$ variables. By indicating the transportation cost matrix $C(\widehat{X})$, we define its components as follows,
@@
s.t. $P\mathbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},\ P\geq\mathbf{0},\ AP^{\top}\mathbf{1}_{n}\geq\mathbf{0}\ .$
@@
s.t. $P\mathbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},\ P\geq\mathbf{0},\ AP^{\top}\mathbf{1}_{n}\geq\mathbf{0}$
@@ 4 Majority Minimization for Solving the Reformulated Problem
Majority minimization aims at solving nonconvex optimization problems, and refers to the process of defining a convex surrogate function that upper bounds the nonconvex objective function, so that optimizing the surrogate function improves the objective function [56, 39]. As the algorithm proceeds, the surrogate function also updates accordingly, which ensures the value of the original objective function keeps decreasing. Following this framework, we define the surrogate function $g(\cdot;\widehat{X}^{k})$ as follows for the $k$-th iterate $\widehat{X}^{k}\in{\cal X}^{m}$:
@@ 4.1 Updating the Surrogate Function (Line 4)
To update the surrogate function, we need to solve problem (8), which is a large-scale linear program. Rather than solving the computationally prohibitive dual problem we solve a lower-dimensional dual problem by using a variant of the FairWASP algorithm proposed by [77]. We adapt FairWASP for cases where $m\neq n$ to find the solution of (8) via applying the cutting plane methods on the Lagrangian dual problems with reduced dimension [^3] As opposed to the scenario where $m=n$, which was tackled by [77].. We choose FairWASP over established commercial solvers due to its computational complexity being lower than other state of the art approaches such as interior-point or simplex method; see Lemma A.2 in Appendix A for more details.